OpenStudy (anonymous):

how to solve x^2-6x+10

OpenStudy (anonymous):

x^2-6x+10=0

OpenStudy (dangerousjesse):

Move everything to the left hand side. Subtract 6 x from both sides: \(x^2-6 x+10 = 0\) Solve the quadratic equation by completing the square. Subtract 10 from both sides: \(x^2-6 x = -10\) Take one half of the coefficient of x and square it, then add it to both sides. Add 9 to both sides: \(x^2-6 x+9 = -1\) Factor the left hand side. Write the left hand side as a square: \((x-3)^2 = -1\) Eliminate the exponent on the left hand side. Take the square root of both sides: \(x-3 = i\) or \(x-3 = -i\) Look at the first equation: Solve for x. Add 3 to both sides: \(x = 3+i\) or \(x-3 = -i\) Look at the second equation: Solve for x. Add 3 to both sides...

OpenStudy (dangerousjesse):

Which leaves you with...?

OpenStudy (anonymous):

x=3-i

OpenStudy (anonymous):

could you show me how to do this using the quadratic formula, please?

OpenStudy (anonymous):

I will give it a go :) \[x=\frac{-6\pm\sqrt{6^2-4\times1\times10}}{2\times1}\] \[x=\frac{-6\pm\sqrt{36-40}}{2}\] \[x=\frac{-6\pm\sqrt{-4}}{2}\] \[x=\frac{-6\pm\sqrt{4}\sqrt{-1}}{2}\] \[x=\frac{-6\pm2i}{2}\] \[x=\frac{-6+2i}{2} or \frac{-6-2i}{2}\] \[x=\frac{2(-3+i)}{2} or \frac{2(-3-i)}{2}\] \[x=i-3 \space or -3-i\] Make sense?

OpenStudy (anonymous):

this whole time I put A in the place of B. makes perfect sense now , thank you!(:

OpenStudy (anonymous):

Haha, cool, nice work.